# Majorana edge and end states in planar Josephson junctions

### v2

@@ III.1 Effects of Rashba SOC
In Al/HgTe Josephson junctions (JJs), where Rashba SOC is the dominant and Dresselhaus SOC is negligibly small (i.e., $\beta\approx 0$), the spin-orbit angle $\theta_{so}\approx 0$ and $\lambda\approx\alpha$ [see Eqs. (3) and (4)]. In this scenario, the system exhibits magneto anisotropyy (i.e., its properties depend on the in-plane magnetic field orientation, $\varphi_{B}$), while crystalline anisotropy is absent (i.e., the system properties are independent of the junction’s crystallographic direction, $\theta_{c}$) Pakizer et al. (2021); Pekerten et al. (2022, 2024b). For the calculations presented in this subsection, the in-plane magnetic field was aligned with the junction direction (i.e., $\varphi_{B}=\pi/2$), which, as discussed in Ref. Pakizer et al. (2021), yields the optimal topological gap for this configuration.
@@
where $j=e=1$ ($j=h=-1$) for electron-like (hole-like) states, $E_{Z}=|\mathbf{E}_{Z}|$, and $k_{so}=m^{\ast}\alpha/\hbar^{2}$. This indicates that zero-phase edge-like MSs of junctions with only Rashba SOC and magnetic field along the junction direction appear when the system is in the TS state and the following conditions are fulfilled,
@@
When the second inequality in Eq.(20) is satisfied (as it is for the parameters used in the numerical simulations), the emergence of edge-like MSs at $\phi=0$ in the limit $W_{S}\rightarrow\infty$ is governed by the condition $E_{Z}\geq\Delta_{0}$. However, since the numerical simulations account for superconducting regions of finite width, the results shown in Fig. 2(a) —where zero-phase edge-like MSs appear at Zeeman energies slightly below $\Delta_{0}$ —exhibit a small deviation from the condition $E_{Z}\geq\Delta_{0}$. The impact of finite-size effects on the topological phase diagram of planar JJs has been explored in Ref.Pekerten et al. (2024b).
@@
In phase-unbiased junctions, the system’s state evolves according to the trajectory of the ground-state phase [see green solid line in Fig. 2(a)] as the Zeeman field is varied. During the $0-\pi$ ground-state jump at $E_{Z}\approx 0.14$ meV, the junction undergoes a transition from the trivial to a TS phase with end-like MSs. A self-tuning mechanism, where edge-like MSs transition into end-like MSs (and/or vice versa) as the Zeeman field varies, appears impractical in phase-unbiased junctions with only Rashba SOC. However, a topologically protected transition in which end- and edge-like MSs can be transformed into each other without exiting the TS state seems feasible by tuning the magnetic flux in phase-biased Josephson junctions (JJs). Since the system remains in the same TS state the topological gap does not close during the process, providing topological protection to the transition. This is illustrated in Fig. 2(a), where an edge-like MS (cyan triangle) can evolve into an end-like MS (magenta dot) by adjusting the Zeeman field (at the value indicated by the vertical dashed line) and varying the phase from 00 to $\pi$.
@@
To visualize the spatial extension of the edge-like MSs along the sample edges perpendicular to the junction, Fig. 2(b) shows the probability density (normalized to its maximum value) of the edge-like MS corresponding to the cyan triangle in Fig. 2(a) at $\phi=0$. The probability density exhibits an oscillatory behavior with multiple maxima of comparable amplitude. Note that the edge-like MSs consistently spread along the entire edge extension, regardless of the width $W_{S}$ of the S regions (see Appendix C). In contrast, the probability density of the end-like MSs corresponding to the magenta square and magenta dot in Fig. 2(a) [see Figs. 2(c) and (d), respectively] is localized at the opposite ends of the junction with wave functions that decay both along and perpendicular to the junction’s direction.
@@
The energy spectrum as a function of $E_{Z}$ is shown in Figs. 2(e)-(g) for $\phi=0,\pi/2,\pi$, respectively. The red solid and blue dotted lines represent the states with energies closest to zero, which evolve into MS as the junction transitions into the TS phase. The vertical lines mark the Zeeman energy boundaries of the first topological region, characterized by the $\mathbb{Z}_{2}$ invariant of the D class, where the topological charge is $Q=-1$ [see Eq. (7)]. Within this region, the absolute value of the $\mathbb{Z}$ invariant of the BDI class equals 1, indicating the presence of a single pair of MSs, as shown by the red lines in Figs. 2(e)-(g). Additionally, regions where a second pair of MSs emerge (blue dashed lines) can also be observed. In these regions, the system remains in the BDI class and supports two pairs of MSs, corresponding to an absolute value of 2 for the $\mathbb{Z}$ invariant of the BDI class. However, it is important to note that the topological gap in regions with multiple pairs of MSs is significantly smaller than in regions where only a single pair of MSs exists [this is particularly clear in Fig. 2(g)]. For this reason, our analysis of edge-like and end-like MSs is focused on topological regions containing only a single pair of MSs.
@@
Figure 3: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb JJ, where the Rashba SOC has been tuned to a negligibly small value and only Dresselhaus SOC is relevant ($\theta_{so}=\pi/2$). The junction and magnetic field orientations are set to $\theta_{c}=0=\varphi_{B}=0$. The green solid line represents the path of the ground-state phase ($\phi_{GS}$) as the Zeeman energy is varied. The vertical dashed line marks a possible transition between a TS state supporting a zero-phase edge-like MS (cyan triangle) and one supporting an end-like Majorana state (magenta dot) during which $E_{Z}$ is kept constant, while $\phi$ is tuned. (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), cyan square (edge-like MS), and magenta dot (end-like MS), respectively. (e)-(g) Energy spectra as a function of the Zeeman energy for $\phi=0$, $\phi=\pi/2$, and $\phi=\pi$, respectively. Red-solid and dashed-blue lines represent states that evolve into MSs as $E_{Z}$ is varied. Vertical dashed lines indicate the boundaries of the first topological region in which only a single pair of MSs (red solid lines) exists. The deviation of the edge-like MS energies (red solid lines) from zero in (f) results from the wavefunction overlap between edge-like MSs on opposite edges [see (c)].
@@ III.3 Effects of combined Rashba and Dresselhaus SOCs
In systems like Al/InSb-based junctions, the coexistence of significant Rashba and Dresselhaus SOC generates a two-fold symmetric spin-orbit field. This, combined with the Zeeman interaction, results in nontrivial magneto anisotropic and crystalline anisotropic effects. The relative strength of Rashba ($\alpha$) and Dresselhaus ($\beta$) SOCs can be tuned by adjusting the Rashba SOC via a gate placed on top of the junction Dartiailh et al. (2021). A particularly interesting regime arises when the Rashba and Dresselhaus SOCs are equally strong. Based on the parametrization introduced in Eqs. (3) and (4), the condition $\alpha=\beta$ corresponds to the SOC angle $\theta_{so}=\pi/4$.
Figure 4: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb JJ, with equal Rashba and Dresselhaus SOC strengths ($\theta_{so}=\pi/4$). The junction and magnetic field orientations are set to $\theta_{c}=3\pi/4$ and $\varphi_{B}=\pi/2$, respectively. The green solid line represents the path of the ground-state phase ($\phi_{GS}$) as the Zeeman energy is varied. The vertical (horizontal) dashed line marks a possible transition between a TS state supporting a zero-phase edge-like MS (cyan triangle/dot) and one supporting an end Majorana state (magenta square) during which $E_{Z}$ ($\phi$) is kept constant while $\phi$ ($E_{Z}$) is tuned. A transition between edge-like (e.g., cyan dot) and end-like (e.g., magenta cross) MSs can also be achieved by solely tuning $E_{Z}$, as the value of $\phi$ self-adjusts and follows the path of the ground-state phase (green solid line). (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), magenta square (end-like MS), and cyan dot (edge-like MS), respectively. (e)-(g) Energy spectra as a function of the Zeeman energy for $\phi=0$, $\phi=\pi/2$, and $\phi=3\pi/5$, respectively. Red-solid and dashed-blue lines represent states that evolve into MSs as $E_{Z}$ is varied. Vertical dashed lines indicate the boundaries of the first topological region in which only a single pair of MSs (red solid lines) exists.
@@
As in the cases discussed in the previous subsections, in phase-biased JJs with both Rashba and Dresselhaus SOC transitions between edge and end-like MSs can also be induced by tuning the magnetic field while keeping the phase fixed at an appropriate value [e.g., following the dotted line from the cyan dot to the magenta square in Fig. 4(a)] or by fixing the magnetic field and tuning the phase difference [e.g, following the dotted line from the cyan triangle to the magenta square in Fig. 4(a)]. Remarkably, the coexistence of Rashba and Dresselhaus SOC allows for transitions between edge and end-like MSs in phase-unbiased junctions, something not observed in systems with only Rashba or only Dresselhaus SOC. As shown in Fig. 4(a), the self-tuning of the ground-state phase (green solid line) as the Zeeman energy varies creates a transition pathway between (red) regions hosting edge-like MSs and (blue) regions hosting end-like MSs. For instance, along the ground-state trajectory, the edge-like MSs at the cyan dot can transition into the end-like MSs at the magenta cross.
@@
Figure 5: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb JJ, with Rashba SOC strength about 2.4 times greater than the Dresselhaus SOC strength ($\theta_{so}=\pi/8$). The junction and magnetic field orientations are set to $\theta_{c}=3\pi/4$ and $\varphi_{B}=\pi/2$, respectively. (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), magenta square (end-like MS), and magenta dot (end-like MS), respectively. (e) Energy spectrum along the path indicated by the vertical dashed line in (a), where $E_{Z}=0.23$ meV and $\phi$ is varied from 0 to $2\pi$. The symbols in (e) indicated the energy of the MSs whose probability densities are plotted in (b)-(d).
@@
Figure 6: (a) Energy spectrum as a function of $W_{S}$ for $\phi=0$ and $E_{Z}=0.25$ meV. The other system parameters were taken as in Fig. 4. Red-solid and dashed-blue lines represent states that evolve into MSs as $W_{S}$ is varied. Vertical dashed lines indicate $W_{S}$-values of 0.4 $\mu$m, 0.8 $\mu$m, 1.2 $\mu$m, and 2 $\mu$m. (b)-(f) Probability density (normalized to its maximum value) for JJs with (b) $W_{S}=0.8\;\mu$m, (c) and (d) $W_{S}=1.2\;\mu$m, and (e) and (f) $W_{S}=2\;\mu$m. The edge-like Majorana pair in (b) evolves into the one shown in (c) (when $W_{S}$ increases from $W_{S}=0.8\;\mu$m to $W_{S}=1.2\;\mu$m) and then into the edge-like pair shown in (e) when $W_{S}=2\;\mu$m. A second Majorana pair [shown in (d)] exists when $W_{S}=1.2\;\mu$m and eventually evolves into the end-like MS displayed in (f).
@@
The probability density (normalized to its maximum value) of the edge-like MSs in a JJ with $W_{S}=0.4\;\mu$m [see Fig. 4(b)] transforms into the edge-like MSs shown in Figs. 6(b), (c), and (e) as $W_{S}$ increases to 0.8 $\mu$m, 1.2 $\mu$m, and 2 $\mu$m, respectively. A second Majorana pair appears at $W_{S}=1.2\;\mu$m [see Fig. 6(d)] and eventually transitions into the end-like Majorana pair displayed in Fig. 6(f) as the extent of the S regions increase to $2\;\mu$m. Although not depicted in Fig. 6, a third pair of MSs exists at $W_{S}=2\;\mu$m.
@@
Due to their distinct localization characteristics, end-like and edge-like MSs can yield different experimental outcomes. Understanding the behavior and properties of these states is, therefore, crucial for correctly interpreting their experimental detection. For instance, the sharp contrast in the zero-bias conductance peak (ZBCP) observed in Josephson junctions (JJs) when the superconducting phase is changed from $\phi=0$ to $\phi=\pi$ Hart et al. (2014) could be related to a transition between edge-like and end-like MSs. At $\phi=0$, edge-like MSs exhibit reduced probability density beneath the contact, leading to a smaller ZBCP amplitude. Conversely, at $\phi=\pi$, end-like MSs have a larger probability density below the contact, resulting in a more prominent ZBCP amplitude. Moreover, because edge-like MSs extend along the entire junction edges, they could act as global interconnects, enabling coupling between distant MSs. This coupling can be controlled by switching between edge-like and end-like MSs, offering a potential mechanism for modulating interactions in topological quantum devices.
Figure 7: Probability density (normalized to its maximum value) for JJs with $W_{S}=1.2\;\mu$m, $W_{N}=0.2\;\mu$m, $E_{z}=0.25$ meV, and chemical potential (a) $\mu_{S(N)}=4$ meV, (b) $\mu_{S(N)}=6$ meV, (c) $\mu_{S(N)}=12$ meV, and (d) $\mu_{S(N)}=15$ meV. The other parameters were taken as in Fig. 4(a), namely, $\theta_{so}=\pi/4$, $\theta_{c}=3\pi/4$, $\phi=0$, and $\varphi_{B}=\pi/2$.

### v3

@@ III.1 Effects of Rashba SOC
In Al/HgTe Josephson junctions (JJs), where Rashba SOC is the dominant and Dresselhaus SOC is negligibly small (i.e., $\beta\approx 0$), the spin-orbit angle $\theta_{so}\approx 0$ and $\lambda\approx\alpha$ [see Eqs. (3) and (4)]. In this scenario, the system exhibits magneto anisotropyy (i.e., its properties depend on the in-plane magnetic field orientation, $\varphi_{B}$), while crystalline anisotropy is absent (i.e., the system properties are independent of the junction’s crystallographic direction, $\theta_{c}$) CATSJJ; Pekerten2022:PRB; Pekerten2024:APL. For the calculations presented in this subsection, the in-plane magnetic field was aligned with the junction direction (i.e., $\varphi_{B}=\pi/2$), which, as discussed in Ref. CATSJJ, yields the optimal topological gap for this configuration.
@@
where $j=e=1$ (j=h=1j=h=-1) for electron-like (hole-like) states, $E_{Z}=|\mathbf{E}_{Z}|$, and $k_{so}=m^{\ast}\alpha/\hbar^{2}$. This indicates that zero-phase propagating edge-like MSs of junctions with only Rashba SOC and magnetic field along the junction direction appear when the system is in the TS state and the following conditions are fulfilled,
@@
When the second inequality in Eq.(20) is satisfied (as it is for the parameters used in the numerical simulations), the emergence of edge-like MSs at $\phi=0$ in the limit $W_{S}\rightarrow\infty$ is governed by the condition $E_{Z}\geq\Delta_{0}$. However, since the numerical simulations account for superconducting regions of finite width, the results shown in Fig. 2(a) —where zero-phase edge-like MSs appear at Zeeman energies slightly below $\Delta_{0}$ —exhibit a small deviation from the condition $E_{Z}\geq\Delta_{0}$. The impact of finite-size effects on the topological phase diagram of planar JJs has been explored in Ref.Pekerten2024:APL.
@@
In phase-unbiased junctions, the system’s state evolves according to the trajectory of the ground-state phase [see green solid line in Fig. 2(a)] as the Zeeman field is varied. During the $0-\pi$ ground-state jump at $E_{Z}\approx 0.14$ meV, the junction undergoes a transition from the trivial to a TS phase with end-like MSs. A self-tuning mechanism, where edge-like MSs transition into end-like MSs (and/or vice versa) as the Zeeman field varies, appears impractical in phase-unbiased junctions with only Rashba SOC. However, a topologically protected transition in which end- and edge-like MSs can be transformed into each other without exiting the TS state seems feasible by tuning the magnetic flux in phase-biased Josephson junctions (JJs). Since the system remains in the same TS state the topological gap does not close during the process, providing topological protection to the transition. This is illustrated in Fig. 2(a), where an edge-like MS (cyan triangle) can evolve into an end-like MS (magenta dot) by adjusting the Zeeman field (at the value indicated by the vertical dashed line) and varying the phase from 0 to $\pi$.
@@
To visualize the spatial extension of the edge-like MSs along the sample edges perpendicular to the junction, Fig. 2(b) shows the probability density (normalized to its maximum value) of the edge-like MS corresponding to the cyan triangle in Fig. 2(a) at $\phi=0$. The probability density exhibits an oscillatory behavior with multiple maxima of comparable amplitude. Note that the edge-like MSs consistently spread along the entire edge extension (see Appendix C). In contrast, the probability density of the end-like MSs corresponding to the magenta square and magenta dot in Fig. 2(a) [see Figs. 2(c) and (d), respectively] is localized at the opposite ends of the junction with wave functions that decay both along and perpendicular to the junction’s direction.
@@
The energy spectrum as a function of $E_{Z}$ is shown in Figs. 2(e)-(g) for $\phi=0,\pi/2,\pi$, respectively. The red solid and blue dotted lines represent the states with energies closest to zero, which evolve into MS as the junction transitions into the TS phase. The vertical lines mark the Zeeman energy boundaries of the first topological region, characterized by the $\mathbb{Z}_{2}$ invariant of the D class, where the topological charge is Q=1Q=-1 [see Eq. (7)]. Within this region, the absolute value of the $\mathbb{Z}$ invariant of the BDI class equals 1, indicating the presence of a single pair of MSs, as shown by the red lines in Figs. 2(e)-(g). Additionally, regions where a second pair of MSs emerge (blue dashed lines) can also be observed. In these regions, the system remains in the BDI class and supports two pairs of MSs, corresponding to an absolute value of 2 for the $\mathbb{Z}$ invariant of the BDI class. However, it is important to note that the topological gap in regions with multiple pairs of MSs is significantly smaller than in regions where only a single pair of MSs exists [this is particularly clear in Fig. 2(g)]. For this reason, our analysis of edge-like and end-like MSs is focused on topological regions containing only a single pair of MSs.
@@
Figure 3: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb $J$, where the Rashba SOC has been tuned to a negligibly small value and only Dresselhaus SOC is relevant ($\theta_{so}=\pi/2$). The junction and magnetic field orientations are set to $\theta_{c}=0=\varphi_{B}=0$. The green solid line represents the path of the ground-state phase ($\phi_{GS}$) as the Zeeman energy is varied. The vertical dashed line marks a possible transition between a TS state supporting a zero-phase edge-like MS (cyan triangle) and one supporting an end-like Majorana state (magenta dot) during which $E_{Z}$ is kept constant, while $\phi$ is tuned. (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), cyan square (edge-like MS), and magenta dot (end-like MS), respectively. (e)-(g) Energy spectra as a function of the Zeeman energy for $\phi=0$, $\phi=\pi/2$, and $\phi=\pi$, respectively. Red-solid and dashed-blue lines represent states that evolve into MSs as $E_{Z}$ is varied. Vertical dashed lines indicate the boundaries of the first topological region in which only a single pair of MSs (red solid lines) exists. The deviation of the edge-like MS energies (red solid lines) from zero in (f) results from the wavefunction overlap between edge-like MSs on opposite edges [see (c)].
@@ III.3 Effects of combined Rashba and Dresselhaus SOCs
In systems like Al/InSb-based junctions, the coexistence of significant Rashba and Dresselhaus SOC generates a two-fold symmetric spin-orbit field. This, combined with the Zeeman interaction, results in nontrivial magneto anisotropic and crystalline anisotropic effects. The relative strength of Rashba ($\alpha$) and Dresselhaus ($\beta$) SOCs can be tuned by adjusting the Rashba SOC via a gate placed on top of the junction PhysRevLett.126.036802. A particularly interesting regime arises when the Rashba and Dresselhaus SOCs are equally strong. Based on the parametrization introduced in Eqs. (3) and (4), the condition $\alpha=\beta$ corresponds to the SOC angle $\theta_{so}=\pi/4$.
Figure 4: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb $J$, with equal Rashba and Dresselhaus SOC strengths ($\theta_{so}=\pi/4$). The junction and magnetic field orientations are set to $\theta_{c}=3\pi/4$ and $\varphi_{B}=\pi/2$, respectively. The green solid line represents the path of the ground-state phase ($\phi_{GS}$) as the Zeeman energy is varied. The vertical (horizontal) dashed line marks a possible transition between a TS state supporting a zero-phase edge-like MS (cyan triangle/dot) and one supporting an end Majorana state (magenta square) during which $E_{Z}$ ($\phi$) is kept constant while $\phi$ ($E_{Z}$) is tuned. A transition between edge-like (e.g., cyan dot) and end-like (e.g., magenta cross) MSs can also be achieved by solely tuning $E_{Z}$, as the value of $\phi$ self-adjusts and follows the path of the ground-state phase (green solid line). (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), magenta square (end-like MS), and cyan dot (edge-like MS), respectively. (e)-(g) Energy spectra as a function of the Zeeman energy for $\phi=0$, $\phi=\pi/2$, and $\phi=3\pi/5$, respectively. Red-solid and dashed-blue lines represent states that evolve into MSs as $E_{Z}$ is varied. Vertical dashed lines indicate the boundaries of the first topological region in which only a single pair of MSs (red solid lines) exists.
@@
[…]
@@
Figure 5: (a) Topological gap character ($\tilde{\Delta}$) as a function of the Zeeman energy $E_{Z}$ and the superconducting phase difference ($\phi$) across an Al/InSb $J$, with Rashba SOC strength about 2.4 times greater than the Dresselhaus SOC strength ($\theta_{so}=\pi/8$). The junction and magnetic field orientations are set to $\theta_{c}=3\pi/4$ and $\varphi_{B}=\pi/2$, respectively. (b)-(d) Probability density (normalized to its maximum value) of the MSs corresponding to the $E_{Z}$ and $\phi$ values marked in (a) by the cyan triangle (edge-like MS), magenta square (end-like MS), and magenta dot (end-like MS), respectively. (e) Energy spectrum along the path indicated by the vertical dashed line in (a), where $E_{Z}=0.23$ meV and $\phi$ is varied from 0 to $2\pi$. The symbols in (e) indicated the energy of the MSs whose probability densities are plotted in (b)-(d).
@@
[…]
@@
Due to their distinct localization characteristics, end-like and edge-like MSs can yield different experimental outcomes. Understanding the behavior and properties of these states is, therefore, crucial for correctly interpreting their experimental detection. For instance, the sharp contrast in the zero-bias conductance peak (ZBCP) observed in Josephson junctions (JJs) when the superconducting phase is changed from $\phi=0$ to $\phi=\pi$ hart2014induced could be related to a transition between edge-like and end-like MSs. At $\phi=0$, edge-like MSs exhibit reduced probability density beneath the contact, leading to a smaller ZBCP amplitude. Conversely, at $\phi=\pi$, end-like MSs have a larger probability density below the contact, resulting in a more prominent ZBCP amplitude. Moreover, because edge-like MSs extend along the junction edges, they could act as global interconnects, enabling coupling between distant MSs. This coupling can be controlled by switching between edge-like and end-like MSs, offering a mechanism for modulating interactions in topological quantum devices and potentially enabling braiding and fusion operations. Figure 8 illustrates how transitions to edge-like Majorana states (MSs) can be employed to interconnect MSs formed in adjacent planar Josephson junctions.
[…]
Figure 8: Schematics of four planar JJs coupled via quantum dots (QDs). Each junction hosts two MSs (red regions), which can be used to encode one topological qubit per junction. The effective coupling between nearby MSs is controlled by QDs QD3 and QD4. Transitions from distant end-like to edge-like MSs enable interconnections between MSs formed in adjacent $J$s (e.g., between $\gamma_{5}$ and $\gamma_{7}$ or between $\gamma_{6}$ and $\gamma_{8}$), where QDs QD1, QD2, QD5, and QD6 can be used to modulate the couplings between MSs along the horizontal direction of the array. Braiding operations can be implemented by appropriately tuning the effective couplings between MSs belonging to different pairs Flensberg2011:PRL; vanHeck2012:NJP.
[…]
III.4 Effects of electrostatic disorder on end-like and edge-like Majorana states
Randomly distributed impurities in semiconductors introduce fluctuations in the electrostatic landscape of proximitized Josephson junctions (JJs), thereby influencing key properties of the topological superconducting state such as the topological gap and the localization of MSs Haim2019:PRL; Paudel2205:PReApplied; sharma2024:arxiv; Laubscher2025:PRB; Escribano2025:PRB. To capture the impact of electrostatic disorder on MSs, we add to the system Hamiltonian [Eq.(1)] an effective disorder potential arising from $N_{imp}$ impurities randomly distributed across the lattice. Following the model described in Ref. Paudel2205:PReApplied, each impurity contributes a local potential with a randomly assigned sign and an amplitude that decays exponentially with distance from the impurity’s position. As a result, the total disorder potential at a given lattice site $\mathbf{r}_{i}=(x_{i},y_{i})$ is determined by the cumulative effect of all nearby impurities. The effective disorder potential at the lattice point $\mathbf{r}_{i}=(x_{i},y_{i})$ is determined by Paudel2205:PReApplied,
$V_{imp}(\mathbf{r}_{i})=\frac{V_{0}\left[f(\mathbf{r}_{i})-\langle f(\mathbf{r}_{i})\rangle\right]}{\sigma_{f}},$ (21)
where $V_{0}$ is the overall potential amplitude,
$f(\mathbf{r}_{i})=\sum_{j=1}^{N_{imp}}(-1)^{j}\exp\left(-\frac{|\mathbf{r}_{i}-\mathbf{R}_{j}|}{\lambda_{imp}}\right),$ (22)
and $\sigma_{f}$ the standard deviation of $f(\mathbf{r}_{i})$. Here $\mathbf{R}_{j}$ denotes the position of the $j$th impurity and $\lambda_{imp}$ characterizes the decay length of the local impurity potential. The potential defined in Eq. (21) satisfies the constraints $\langle V_{imp}\rangle=0$ and $\langle V_{imp}^{2}\rangle=V_{0}^{2}$.
A random sample of $N_{imp}=3171$ impurities in a $(2.1\;\mu{\rm m}\times 5\;\mu{\rm m})$ JJ, corresponding to an impurity concentration of 3% ($\sim 300\;\mu{\rm m}^{-2}$), and $\lambda_{imp}=70$ nm were used in the numerical simulations. The position dependence of the impurity potential normalized to its maximum amplitude is shown in Fig. 9(b), where the tiny black dots mark the locations of the randomly distributed impurities.
Figure 9(a) displays the difference $\Delta E$ between the two lowest positive-energy states (which defines the topological gap in the topological superconducting phase) as a function of the disorder potential amplitude $V_{0}$. Results are shown for both end-like (squares) and edge-like (dots) MSs. Overall, $\Delta E$ decreases with increasing $V_{0}$, indicating a deterioration of the topological states as disorder becomes stronger. However, this decay is not strictly monotonic and $\Delta E$ exhibits some local maxima, a feature that appears generic. Simulations for multiple disorder realizations (not shown) reveal the same qualitative nonmonotonic behavior, though the precise locations of these maxima vary with the specific disorder configuration.
Interestingly, weak disorder can enhance the topological gap of end-like MSs compared to the clean case. This can be seen in Fig. 9(a), where the topological gap protecting the end-like MSs is enhanced with respect to the clean cases for finite $V_{0}$ values up to 0.4 meV. This observation is consistent with previous studies, which predicted that introducing weak disorder in JJs can enhance the robustness of the topological superconducting phase Haim2019:PRL. The topological gap for edge-like MSs tends to decrease more rapidly with $V_{0}$ in the weak-disorder regime but becomes larger than that for end-like MSs at higher $V_{0}$. This crossover may not be universal and could shift when the Zeeman field is varied. For clarity, the Zeeman field values were fixed so that the topological gap of end-like and edge-like MSs is the same in the absence of impurities. Nevertheless, at each $V_{0}$, the gap could likely be further optimized by tuning $E_{Z}$.
The probability densities of MSs for the $V_{0}$ values indicated by the vertical dashed lines in Fig. 9 (a) (namely, $V_{0}=0.2,0.8,1.0$ and 1.8 meV) are displayed in Figs. 9(c)-(j), where the left and right panels correspond to the end-like and edge-like states, respectively. In the weak-disorder regime, end-like (edge-like) MSs remain well localized near the system ends (edges). However, as the disorder strength increases, the MSs gradually change their localization regions, penetrating deeper into the bulk and eventually approach each other (a similar trend was reported in Ref. Paudel2205:PReApplied for the case of end-like MSs). This behavior poses additional challenges for the practical use of MSs in planar JJs for quantum computing, as maintaining disorder in the weak regime is essential to prevent uncertainties in the MS positions and their reliable manipulation.